The arrow strikes the target with a speed of 45 m / s at an angle of −25 ° with respect to the horizontal. Calculate the horizontal and vertical components of the arrow's final velocity.
1
Expert's answer
2020-12-06T17:24:57-0500


The x-component of the velocity vector "\mathbf{v}" is:


"v_x = 45\cdot \cos 25\degree \approx 40.78\space m/s"

The y-component of the velocity vector "\mathbf{v}" is:


"v_y = -45\cdot \sin 25\degree \approx -19.02\space m/s"

Answer. "v_x = 40.78m\/s," "v_y = -19.02m\/s".